OpenStudy (anonymous):

3(x-5)=7(2x+3)-4(x-2)-5(x-2)+2

OpenStudy (anonymous):

Wanna work this one out with me, Archana? I can walk you through it.

OpenStudy (anonymous):

Sure..

OpenStudy (anonymous):

Awesome. First thing we want to do is get rid of the parentheses. So how do we get rid of the parentheses in the first part, 3(x-5) ?

OpenStudy (anonymous):

3x-15

OpenStudy (anonymous):

Perfect. We have 3x-15 = 7(2x+3)-4(x-2)-5(x-2)+2

OpenStudy (anonymous):

Could you tell me what we have after continuing with the right side? Be careful with negative numbers.

OpenStudy (anonymous):

14x+21-4x+8-5x+10+2

OpenStudy (anonymous):

Awesome! 3x-15 = 14x+21-4x+8-5x+10+2

OpenStudy (anonymous):

Now what do we do?

OpenStudy (anonymous):

Hint: All the work we do will be on the right side of the = :D

OpenStudy (anonymous):

Insert like terms together..

OpenStudy (anonymous):

Yup! So what do we have?

OpenStudy (anonymous):

3x-15=14x-4x-5x+21+8+10+2

OpenStudy (anonymous):

Excellent. Now just add up the similar terms and lemme know what we have.

OpenStudy (anonymous):

3x -15 = 5x + 41 is what I got. Sound good?

OpenStudy (anonymous):

3x-15=5x+41

OpenStudy (anonymous):

Sweetness! And what do we do now?

OpenStudy (anonymous):

That's rite I was a bit late..

OpenStudy (anonymous):

Don't worry, you got it really quickly. I just had a calculator with me. :P

OpenStudy (anonymous):

Transfer left elements to right, is dat correct..

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Is the answer 28

OpenStudy (anonymous):

well, you'll want x on one side and constants (non-variables) on another side. Which side do you want X to be?

OpenStudy (anonymous):

Any wr..

OpenStudy (anonymous):

To be honest, I think the answer will be negative. Let's keep x on the right since there's 'more Xs' over there.

OpenStudy (anonymous):

Ok..

OpenStudy (anonymous):

how can we get rid of the 3x on the left?

OpenStudy (anonymous):

3x-15=5x+41 Here's the problem again so you don't have to keep scrolling.

OpenStudy (anonymous):

Is it transferring to the rite side,,

OpenStudy (anonymous):

Yup. But how do we move it there?

OpenStudy (anonymous):

5x-3x=41+15

OpenStudy (anonymous):

-2x=56

OpenStudy (anonymous):

x=56/2

OpenStudy (anonymous):

Very close.

OpenStudy (anonymous):

Sorry x=56/-2

OpenStudy (anonymous):

Ah, yes. That is right.

OpenStudy (anonymous):

x=-28

OpenStudy (anonymous):

That's right :D But just one thing:

OpenStudy (anonymous):

Hurray, cheers..

OpenStudy (anonymous):

Wt is dat

OpenStudy (anonymous):

When you did: 5x-3x=41+15 There was a typo. As this would have given you 2x = 56

OpenStudy (anonymous):

Should have been: 3x-15=5x+41 into -41 + -15 = 5x-3x

OpenStudy (anonymous):

But you ended up compensating later, so -28 is correct. ^_^

OpenStudy (anonymous):

Thank's for your help, wt is ur name..

OpenStudy (anonymous):

Real name is Mohammed. But I go by uber on the internet. ;)

OpenStudy (anonymous):

Gr8 nice soling the problem with u k bye take care..

OpenStudy (anonymous):

You're welcome, enjoy!

OpenStudy (anonymous):

Which place r u from..

OpenStudy (anonymous):

R u a teacher?

OpenStudy (anonymous):

I'm afghan, but lived in the US pretty much my whole life. I'm just a college calculus student helpin' out others.